OpenStudy (anonymous):

The perimeter, p, of a rectangular field can be found using the formula p = 2l + 2w where l is the length of the field and w is the width. Which equation could be used to solve for l? A: l = B: l = C: l = p − 2w D: l = p − w

jimthompson5910 (jim_thompson5910):

Solving for L gives us... \[\Large P = 2L + 2W\] \[\Large P-2W = 2L\] \[\Large 2L = P-2W\] \[\Large L = \frac{P-2W}{2}\]